Question:Solve the equation find the value of y for each valueof x??
y+3x=8;x=-2,0,2

Answer by nyc_function(2741)

y + 3x = 8; x = -2, 0, 2


y = -3x + 8


Let x = -2


y = -3(-2) + 8


y = 6 + 8


y = 14


==============


Let x = 0


y = -3x + 8


y = -3(0) + 8


y = 0 + 8


y = 8


=============


Let x = 2

y = -3x + 8

y = -3(2) + 8

y = -6 + 8

y = 2